OpenStudy (anonymous):

I have another one I am stumped on slope -2/3, y-intercept (0,5)

OpenStudy (anonymous):

leave in slope intercept form

OpenStudy (anonymous):

okay, first you need to put it in point slope form [ y1-y2=m(x1-x2) ] so plug in the numbers and get y-5=-2/3(x-0) from there, you distribute the -2/3 inside the parenthesis y-5=-2/3x then simplify- y=-2/3x+5 ---- this is slope intercept form VOILA, MATH! cheers, and let me know if you have any more troubles!

OpenStudy (anonymous):

Yay thank you I am having so much trouble with this for some reason